Circular No. 8864

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2007cw–2007gv Further to IAUC 8860, various CBETs have announced numerous new

supernovae discovered on unfiltered CCD frames: 2007fb, 2007fc, 2007fo, 2007fq, 2007fr, 2007fs, 2007ge, 2007go, and 2007gr by W. Li et al. (LOSS); 2007fp by T.-w. Chen et al. (Taiwan Supernova Survey; cf. IAUC 8843); 2007ft by A. Klotz (cf. IAUC 8860); 2007gi by K. Itagaki (cf. IAUC 8843); 2007gj and 2007gq by L. A. G. Monard (cf. IAUC 8843); 2007gk by A. Sehgal, J. Newton, and T. Orff (cf. IAUC 8851); and the rest by the “Nearby Supernova Factory” collaboration (on CBETs 990, 991, 996, 1007, 1009, 1011, 1013, 1014, 1024, 1026, 1031, and 1040). Discovery observations tabulated for those objects reported to be brighter than mag 17:

SN  2007 UT  α_{2000}  δ_{2000}  Mag.  Offset
2007fb July 3.47 23^{h}56^{m}52.^{s}37 + 5^{o}30′31′′.8 15.7 12′′.8 W, 2′′.5 N
2007fc July 8.45 23 57 33.89 −21 59 53.0 16.9 22′′.9 E, 8′′.5 N
2007fp Feb. 26.79 10 42 18.21 − 0 22 38.4 16.3 3′′.6 E, 1′′.6 S
2007fs July 15.44 22 01 40.44 −21 30 29.6 15.4 12′′.4 E, 2′′.5 N
2007ft July 15.02 12 47 03.7 −41 35 22 16.0 73′′ E, 21′′ S
2007gi July 31.49 12 01 23.42 +61 53 33.8 16.3 23′′ W, 11′′ S
2007gj Aug. 4.12 2 19 33.85 −37 48 44.2 16.2 44′′ W, 22′′ N
2007gq Aug. 11.70 12 35 11.22 − 0 13 21.6 16.6 8′′ E, 6′′ S
2007gr Aug. 15.51 2 43 27.98 +37 20 44.7 13.8 24′′.8 W, 15′′.8 N

Published data on CBETs: 2007fq in MCG –04-48-19, CBET 1001; 2007gj in ESO 298-28, CBET 1022; 2007gq in NGC 4541, CBET 1033. Data for type-Ia supernovae: 2007fb in UGC 12859, CBETs 992, 993; 2007fc in ESO 538-G8, CBETs 995, 998; 2007fr, CBET 1001; 2007fs in ESO 601-G5, CBETs 1002, 1003 (the offset above corrected from that given on CBET 1002); 2007ft in NGC 4677, CBETs 1004, 1005; 2007ge in MCG +08-1-13, CBETs 1012, 1014; 2007gi in NGC 4036, CBETs 1017, 1021; 2007gk in MCG +11-20-27, CBETs 1023, 1025. Data for the type-Ib or -Ic supernovae: 2007fo in NGC 7714, CBETs 997, 1001; 2007go in ESO 475-G16, CBETs 1028, 1037; 2007gr in NGC 1058, CBETs 1034, 1036, 1041. 2007fp in NGC 3340 is a type-II supernova (cf. CBETs 976 and 1000). SNe 2007db, 2007dy, 2007ea, 2007eb, 2007eq, 2007ff, 2007fj, 2007fn, 2007gg, and 2007gl are also of type Ib or Ic; 2007dp, 2007dq, 2007dw, 2007ed, 2007ef, 2007eh, 2007el, 2007em, 2007er–2007eu, 2007ew, 2007fa, 2007fe, 2007fg, 2007fk– 2007fm, 2007fv, 2007fw, 2007fy–2007ga, 2007gh, 2007gm, and 2007gs are of also of type II (the remainder are type-Ia supernovae).

2007 August 24  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2007 CBAT  Daniel W. E. Green